40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Mar 13 03:12:47 2013, Manasa, Jenne, Update, Locking, TWO arms TWO colors 
    Reply  Wed Mar 13 08:34:33 2013, yuta, Update, Locking, TWO arms TWO colors 
       Reply  Wed Mar 13 16:29:25 2013, Manasa, Update, Locking, TWO arms TWO colors 
    Reply  Wed Mar 13 10:26:58 2013, Manasa, Update, Locking, We're still good!! 
    Reply  Mon Mar 25 15:01:33 2013, Manasa, Update, Locking, TT1 drift 
       Reply  Mon Mar 25 15:24:41 2013, rana, Update, Locking, TT1 drift 
Message ID: 8283     Entry time: Wed Mar 13 08:34:33 2013     In reply to: 8282     Reply to this: 8289
Author: yuta 
Type: Update 
Category: Locking 
Subject: TWO arms TWO colors 

- I took the shutter from AS table to use it for the PSL green. It was sitting near MC REFL path unused (elog #8259).

- If X green lock is not tight, maybe temporarily increasing loop gain helps. This can be done by increasing the amplitude of the frequency modulation or increasing green refl PD gain. Also, if X green beam spot is too wiggly compared with Y green, it is maybe because of air flow from the air conditioner (elog #6849). I temporarily turned it off when I did X green steering last summer.

- X green transmission on PSL table reached ~270 uW last summer (elog #6849, elog #6914). Y green transmission is now ~240 uW and ~2700 counts at maximum. So, X green transmission should reach ~3000 in counts.

- Did you have to re-align TRX path? We moved the harmonic separator on X end table horizontally to avoid IR TRX clipping before beam centering on X arm (elog #8162). I wonder what is the current situation after the beam centering.

ELOG V3.1.3-